You push on a box with a force of 5.0 N. the force of the friction on the box pushes back with a force of 2.0N. If the mass of the box is 2.0kg what is the acceleration of the box ?
1
Expert's answer
2013-01-29T07:43:19-0500
The net force on the box(assuming the direction of our pushing as positive) F = 5.0 N - 2.0 N = 3.0 N ;

According to the Newton's 2nd Law, F = m*a;
Acceleration a = F / m = 3.0 N / 2.0 kg = 1.5 m/s^2

Answer: 1.5 m/s^2